OpenStudy (anonymous):

Please Help Me!! (**Mom**), How much is 2/7 of 1 3/4? Is is .50?? Thanks!! :)

OpenStudy (superdavesuper):

Yes u got the right answer :)

OpenStudy (anonymous):

Yes, it is .5 since \(\large 1\frac{3}{4} = \frac{7}{4}\) and 2/7 of this is \(\large \frac{2}{7}\cdot \frac{7}{4} = \frac{1}{2}\).

OpenStudy (anonymous):

Awesome, thanks to you both :)